OpenStudy (anonymous):

Two-thirds of a number increased by one-third is equal to seventeen. Find the number.

OpenStudy (anonymous):

let number=x ite 2/3rd=2x/3 on increasing it by 1/3rd means in it 1/3 part of (2x/3) is added to 2x/3 2x/3+1/3*(2x/3)=17 2x/3+2x=17 (2x+6x)/3=17 8x/3=17 8x=17*3=51 x=51/8